OpenStudy (anonymous):

show that the curves xy = a^2 and x^2 + y^2 = 2a^ touch each other.

OpenStudy (anonymous):

is that \(x^2+y^2=2a^2\)

OpenStudy (anonymous):

anyways observe \(2xy=2a^2\) hence \((x-y)^2=x^2+y^2-2xy=2a^2-2a^2=0\) therefore they meet when \(x=y\)